OpenStudy (anonymous):

What is the sum of an 8–term geometric series if the first term is 13 and the last term is –3,639,168? –3,599,175 –3,359,230 –3,119,285 –2,879,340

ganeshie8 (ganeshie8):

first, find the common ratio r

OpenStudy (anonymous):

Last term of a GP = \[a * r^{n-1}\] => –3639168 = 13*r^7 And that gives r = -6 Now, sum of a GP = \[\frac{a(1-r^n)}{1-r}\]